# A geometric approach toconjugation-invariant random permutations

### v3

@@ 6.1 High records
For each $n$, let $t^{(n)}$ be a cycle type of size $n$ such that $t_{1}^{(n)}=0$, and $\mathcal{P}$ be a geometric construction. Set $C:=[0,1/2]\times[1/2,1]$. Define $Z_{\text{left}}=\left(U_{\text{left}},V_{\text{left}}\right)$ as the leftmost point in $\mathcal{P}$, $Z_{\text{left}}^{C}=\left(U_{\text{left}}^{C},V_{\text{left}}^{C}\right)$ as the leftmost point of $\mathcal{P}\cap C$, and let $Z_{\text{up}}$, $Z_{\text{up}}^{C}$ be the upmost analogues. Those points are well defined w.h.p. as $n\to\infty$ and:
(.$\left(\big{.}nU_{\text{left}},n\left(1-V_{\text{up}}\right)\right)\underset{n\to\infty}{\longrightarrow}\left(E,E^{\prime}\right)\quad;\quad nU_{\text{left}}^{C}\underset{n\to\infty}{\longrightarrow}2E\quad;\quad n\left(1-V_{\text{up}}^{C}\right)\underset{n\to\infty}{\longrightarrow}2E$
@@
This proves the first convergence in distribution. For the second one, use 3.3 and write, conditionally given $\left|\mathcal{P}\cap C\right|$:
@@
According to 3.2, $\left|\mathcal{P}\cap C\right|$ is concentrated around $n/4$. This readily implies $\mathbf{P}\left(nU_{\text{left}}^{C}\geq s\right)\to e^{-s/2}$ as $n\to\infty$, and proves the second convergence in distribution. The third one follows similarly.
@@
Proof of Theorem 2.10.
Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$. Decompose it into its points outside the diagonal $\check{\mathcal{P}}$ and its points on the diagonal $\mathcal{P}_{\Delta}$. We shall use the notation of 6.1, applied to the set $\check{\mathcal{P}}$ of size $\check{n}:=n-t_{1}^{(n)}$. According to this lemma, w.h.p. the point $Z_{\text{left}}$ is above the diagonal $\Delta$ and the rectangle $[U_{\text{left}},U_{\text{left}}^{C}]\times[V_{\text{left}},V_{\text{left}}^{C}]$ does not intersect $\Delta$. Likewise, w.h.p. the point $Z_{\text{up}}$ is above the diagonal and the rectangle $[U_{\text{up}}^{C},U_{\text{up}}]\times[V_{\text{up}}^{C},V_{\text{up}}]$ does not intersect $\Delta$. Under those events, the records of $\mathcal{P}$ are all contained in
@@
This zone is shown on the left-hand side of Figure 5. Note that the rectangles $[U_{\text{left}},U_{\text{left}}^{C}]\times[V_{\text{left}},V_{\text{left}}^{C}]$ and $[U_{\text{up}}^{C},U_{\text{up}}]\times[V_{\text{up}}^{C},V_{\text{up}}]$ may be flat, i.e. the identities $U_{\text{left}}=U_{\text{left}}^{C}$ and $V_{\text{up}}^{C}=V_{\text{up}}$ may occur with non-negligible probability.
Invertly, all points of $\mathcal{P}_{\Delta}$ in $[0,U_{\text{left}}]^{2}\cup[V_{\text{up}},1]^{2}$, as well as the ones which are high records in $\mathcal{P}\cap C$, are high records in $\mathcal{P}$. From the previous two observations we deduce the following inequality w.h.p. as $n\to\infty$:
@@
where $S_{n}:=\left|\mathcal{P}_{\Delta}\cap[0,U_{\text{left}}]^{2}\right|+\left|\mathcal{P}_{\Delta}\cap[V_{\text{up}},1]^{2}\right|$. According to 6.1, $U_{\text{left}}^{C}$ is a $\mathcal{O}_{\mathbf{P}}\left(1/\check{n}\right)$. For any $M>0$, we may apply 3.2 to the rectangle $[0,M/\check{n}]\times[0,1]$ to obtain $\left|\check{\mathcal{P}}\cap\left([0,M/\check{n}]\times[0,1]\right)\right|\leq 2M$ w.h.p. as $n\to\infty$. Since $U_{\text{left}}^{C}\leq M/\check{n}$ with probability tending to $1$ as $M\to\infty$, uniformly in $n$, this implies that the variable $\left|\check{\mathcal{P}}\cap\left([U_{\text{left}},U_{\text{left}}^{C}]\times[V_{\text{left}},V_{\text{left}}^{C}]\right)\right|$ is a $\mathcal{O}_{\mathbf{P}}(1)$. With the same reasoning applied to $1-V_{\text{up}}^{C}$ and the corresponding rectangle, we deduce:
@@
Now it remains to analyse both terms in the sum. According to 3.3 and 3.2, $\check{\mathcal{P}}\cap C$ is a family of i.i.d. uniform points in $C$, with size $\check{n}/{4}+o_{\mathbf{P}}\left(\check{n}\right)$. We can thus apply Theorem 2.9 to deduce
@@
in distribution. Let us turn to the fixed points. $\check{\mathcal{P}}$ and $\mathcal{P}_{\Delta}$ are independent, so $(U_{\text{left}},V_{\text{up}})$ and $\mathcal{P}_{\Delta}$ are as well. Therefore, conditionally given $(U_{\text{left}},V_{\text{up}})$, $S_{n}$ follows a ${\rm Binom}\left(t_{1}^{(n)},U_{\text{left}}\!+\!1\!-\!V_{\text{up}}\right)$ law. Recall from 6.1 that
@@
in distribution. If ${t_{1}^{(n)}}/{\check{n}}=\mathcal{O}(1)$ as $n\to\infty$ then $S_{n}=\mathcal{O}_{\mathbf{P}}(1)$ (use e.g. the Bienaymé–-Chebyshev inequality), whereas if ${t_{1}^{(n)}}/{\check{n}}\to\infty$ as $n\to\infty$ then
@@
in distribution. Now we can use Equations 21, 22 and 23 to identify three regimes in the asymptotics of $\mathrm{rec}^{\uparrow}\left(\mathcal{P}\right)$.
@@
Assume $\frac{\check{n}}{n/\sqrt{\log n}}\to\alpha\in(0,\infty)$. Then $\log\check{n}\sim\log n$ as $n\to\infty$, and since $t_{1}^{(n)}=n-\check{n}$, we deduce $\frac{t_{1}^{(n)}}{\check{n}\sqrt{\log\left(\check{n}\right)}}\to 1/\alpha$. Thus we need to understand the interplay between the limit laws of Equations 22 and 23. To do this, notice that conditionally given $\left(U_{\text{left}},V_{\text{up}}\right)$, the variables $S_{n}$ and $\check{\mathcal{P}}\cap C$ are independent. Recall that under this conditioning, $S_{n}$ follows a ${\rm Binom}\left(t_{1}^{(n)},U_{\text{left}}\!+\!1\!-\!V_{\text{up}}\right)$ law. According to 3.4, the conditional law of ${\rm Perm}\left(\check{\mathcal{P}}\cap C\right)$ given $\left(U_{\text{left}},V_{\text{up}}\right)=\left(U_{(1)},U_{(n)}\right)$ and $\left|\check{\mathcal{P}}\cap C\right|$ is uniform. From this property we deduce independence of the limit laws in Equations 23 and 22, whence:
@@ 6.2 Low records
By 3.2, $\left|\mathcal{P}\cap B_{n}\right|$ is distributed like a sum of three ${\rm Binom}\left(\lfloor n/3\rfloor\pm 1,\left(1-\epsilon/\sqrt{n}\right)\big{.}\epsilon/\sqrt{n}\right)$, and likewise for $B_{n}^{*}$. The first claim readily follows from standard concentration inequalities. Then w.h.p. as $n\to\infty$, the sets $\mathcal{P}\cap B_{n}$ and $\mathcal{P}\cap B_{n}^{*}$ are non-empty and $Z_{\text{down}}^{B}$, $Z_{\text{left}}^{B^{*}}$ are well defined. According to 3.3, conditionally given $\left|\mathcal{P}\cap B_{n}\right|$, the set $\mathcal{P}\cap B_{n}$ consists of i.i.d. uniform points in $B_{n}$. We can thus write, for any $s>0$ and large enough $n$, conditionally given $\left|\mathcal{P}\cap B_{n}\right|$:
@@
Define $C_{n}:=\left[0,\epsilon/\sqrt{\check{n}}\right]^{2}$. Then $\left|\check{\mathcal{P}}\cap C_{n}\right|$ is distributed like a sum of three ${\rm Binom}\left(\lfloor\check{n}/3\rfloor\pm 1,\epsilon^{2}/\check{n}\right)$ variables by 3.2, and $\left|\mathcal{P}_{\Delta}\cap C_{n}\right|$ follows a ${\rm Binom}\left(t_{1}^{(n)},\epsilon/\sqrt{\check{n}}\right)$ law. Therefore there exists $\delta(\epsilon)=\underset{\epsilon\to 0}{o}\left(1\right)$, independent of $n$, such that:
@@
Define $B_{n}:=\left[0,\epsilon/\sqrt{\check{n}}\right]\times\left[\epsilon/\sqrt{\check{n}},1\right]$ and $B_{n}^{*}:=\left[\epsilon/\sqrt{\check{n}},1\right]\times\left[0,\epsilon/\sqrt{\check{n}}\right]$. According to 6.2 we can define w.h.p. $Z_{\text{down}}^{B}=\left(U_{\text{down}}^{B},V_{\text{down}}^{B}\right)$ as the downmost point in $\check{\mathcal{P}}\cap B_{n}$ and $Z_{\text{left}}^{B^{*}}=\left(U_{\text{left}}^{B^{*}},V_{\text{left}}^{B^{*}}\right)$ as the leftmost point in $\check{\mathcal{P}}\cap B_{n}^{*}$. See the right-hand side of Figure 5 for a representation.
@@
|.$\left|\Big{.}\mathrm{rec}^{\downarrow}\left(\mathcal{P}\right)-\mathrm{rec}^{\downarrow}\left(\check{\mathcal{P}}\cap B_{n}\right)-\mathrm{rec}^{\downarrow}\left(\check{\mathcal{P}}\cap B_{n}^{*}\right)\right|\leq\left|\mathcal{P}\cap\left(\left[\epsilon/\sqrt{\check{n}},U_{\text{left}}^{B^{*}}\right]\times\left[\epsilon/\sqrt{\check{n}},V_{\text{down}}^{B}\right]\right)\right|.$ (25)
Using 6.2 and the fact that $t_{1}^{(n)}=\mathcal{O}\left(\sqrt{n}\right)$ we obtain:
@@
as $n\to\infty$, for fixed $\epsilon>0$. Thus it suffices to study the records in $B_{n}$ and $B_{n}^{*}$. We can apply 3.3, 6.2 and Theorem 2.9 to obtain
@@
in probability, and likewise for $B_{n}^{*}$. Using Equations 25 and 24 we deduce that for any $\eta>0$:
@@
in distribution. By 3.2 and since ${\check{n}}/{(t_{1}^{(n)})^{2}}\to 0$, w.h.p. it holds that $\check{\mathcal{P}}\cap[0,U_{\Delta}]^{2}=\emptyset$. Define two bands $B_{\Delta}:=[0,U_{\Delta}]\times[U_{\Delta},1]$ and $B_{\Delta}^{*}:=[U_{\Delta},1]\times[0,U_{\Delta}]$. Then under the event $\left\{\check{\mathcal{P}}\cap[0,U_{\Delta}]^{2}=\emptyset\right\}$ we may write:
@@
Using 3.2 and 3.3 conditionally given $U_{\Delta}$, $\check{\mathcal{P}}\cap B_{\Delta}$ is a family of i.i.d. uniform points in $B_{\Delta}$ whose size is distributed like a sum of three ${\rm Binom}\left(\lfloor\check{n}/3\rfloor\pm 1,U_{\Delta}(1-U_{\Delta})\right)$ variables. The same holds for $B_{\Delta}^{*}$.
  • If $n=\mathcal{O}\left(t_{1}^{(n)}\right)$ then, by Equation 26, $\left|\mathcal{P}\cap B_{\Delta}\right|$ and $\left|\mathcal{P}\cap B_{\Delta}^{*}\right|$ are both $\mathcal{O}_{\mathbf{P}}(1)$ as $n\to\infty$. Thus by Equation 27, $\mathrm{rec}^{\downarrow}\left(\mathcal{P}\right)$ is also a $\mathcal{O}_{\mathbf{P}}(1)$.
  • If $t_{1}^{(n)}=o(n)$ then, by Equation 26, $\frac{t_{1}^{(n)}}{n}\left|\mathcal{P}\cap B_{\Delta}\right|\underset{n\to\infty}{\longrightarrow}{\rm Exp}\left(1\right)$ in distribution. Therefore:
@@
    in probability. In particular $\left|\mathcal{P}\cap B_{\Delta}\right|\to\infty$ in probability and we can apply Theorem 2.9 to deduce:
@@
    in probability. The same holds for $B_{\Delta}^{*}$, and the previous two equations along with Equation 27 yield:
@@
This concludes the proof of Theorem 2.11. ∎
Proof of 2.12.
We use the same notation as in the proof of Theorem 2.11. Recall that under $E_{n}$:
@@
Suppose that $t_{1}^{(n)}+2t_{2}^{(n)}=n$, i.e. that $\tau_{n}$ is a random involution with $t_{1}^{(n)}$ fixed points. Then the sets $\check{\mathcal{P}}\cap B_{n}$ and $\check{\mathcal{P}}\cap B_{n}^{*}$ are a.s. symmetric with respect to the diagonal. Therefore ${\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}\right)$ and ${\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}^{*}\right)$ are a.s. inverse of each other, and $\mathrm{rec}^{\downarrow}\left({\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}\right)\right)=\mathrm{rec}^{\downarrow}\left({\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}^{*}\right)\right)$ almost surely. Using Equations 24, 28 and 29 this yields:
@@
Suppose that $t_{1}^{(n)}+3t_{3}^{(n)}=n$, i.e. that $\tau_{n}$ is a random product of $3$-cycles. We claim the following: conditionally under the event $E_{n}$ and given the family $\mathcal{P}\cap B_{n}$, the permutation ${\rm Perm}\left(\mathcal{P}\cap B_{n}^{*}\right)$ has size $\left|\mathcal{P}\cap B_{n}\right|$ and is uniformly random. Let us explain why.
@@
This holds for each triple $U_{3,k}^{1},U_{3,k}^{2},U_{3,k}^{3}$ in the construction of $\mathcal{P}$. Under the event $E_{n}$, only two possibilities remain: either $Z_{3,k}^{1},Z_{3,k}^{2},Z_{3,k}^{3}\in D_{n}$, or exactly one point is in $B_{n}$ and another one is in $B_{n}^{*}$, and they share a coordinate. The conditional law of $\mathcal{P}\cap B_{n}^{*}$ under $E_{n}$ and $\mathcal{P}\cap B_{n}$ can thus be described as follows: write $\mathcal{P}\cap B_{n}=\{(U_{i},V_{i})\}$; let $(W_{i})$ be i.i.d. ${\rm Unif}\left([\epsilon_{n},1]\right)$ variables; and set $\mathcal{P}\cap B_{n}^{*}=\{(W_{i},U_{i})\}$. One can easily see that for any distinct $u_{1},\dots,u_{r}$, if $W_{1},\dots,W_{r}$ are i.i.d. ${\rm Unif}\left([\epsilon_{n},1]\right)$ variables then ${\rm Perm}\left((W_{1},u_{1}),\dots,(W_{r},u_{r})\right)$ is uniformly random (e.g. switching $W_{i}$ with $W_{j}$ shows that this is the Haar measure on the symmetric group). This proves that conditionally under the event $E_{n}$ and given the family $\mathcal{P}\cap B_{n}$, the permutation ${\rm Perm}\left(\mathcal{P}\cap B_{n}^{*}\right)$ has size $\left|\mathcal{P}\cap B_{n}\right|$ and is uniformly random.
This claim implies that conditionally under the event $E_{n}$ and given the number $\left|\mathcal{P}\cap B_{n}\right|$, the permutations ${\rm Perm}\left(\mathcal{P}\cap B_{n}\right)$ and ${\rm Perm}\left(\mathcal{P}\cap B_{n}^{*}\right)$ have size $\left|\mathcal{P}\cap B_{n}\right|$ and are independent uniform. Using Equations 24, 28 and 29 we deduce:

### v4

@@ 6.1 High records
For each $n$, let $t^{(n)}$ be a cycle type of size $n$ such that $t_{1}^{(n)}=0$, and $\mathcal{P}$ be a geometric construction. Set $C:=[0,1/2]\times[1/2,1]$. Define $Z_{\text{left}}=\left(U_{\text{left}},V_{\text{left}}\right)$ as the leftmost point in $\mathcal{P}$, $Z_{\text{left}}^{C}=\left(U_{\text{left}}^{C},V_{\text{left}}^{C}\right)$ as the leftmost point of $\mathcal{P}\cap C$, and let $Z_{\text{up}}$, $Z_{\text{up}}^{C}$ be the upmost analogues. Those points are well-defined w.h.p. as $n\to\infty$ and:
$\left(\big.nU_{\text{left}},n\left(1-V_{\text{up}}\right)\right)\underset{n\to\infty}{\longrightarrow}\left(E,E^{\prime}\right)\quad;\quad nU_{\text{left}}^{C}\underset{n\to\infty}{\longrightarrow}2E\quad;\quad n\left(1-V_{\text{up}}^{C}\right)\underset{n\to\infty}{\longrightarrow}2E$
@@
This proves the first convergence in distribution. For the second one, use ˜3.3 and write, conditionally given $\left|\mathcal{P}\cap C\right|$:
@@
According to ˜3.2, $\left|\mathcal{P}\cap C\right|$ is concentrated around $n/4$. This readily implies $\mathbf{P}\left(nU_{\text{left}}^{C}\geq s\right)\to e^{-s/2}$ as $n\to\infty$, and proves the second convergence in distribution. The third one follows similarly.
@@
Proof of Theorem˜2.10.
Let $\mathcal{P}$ be a geometric construction of $\tau_{n}$. Decompose it into its points outside the diagonal $\check{\mathcal{P}}$ and its points on the diagonal $\mathcal{P}_{\Delta}$. We shall use the notation of ˜6.1, applied to the set $\check{\mathcal{P}}$ of size $\check{n}:=n-t_{1}^{(n)}$. According to this lemma, w.h.p. the point $Z_{\text{left}}$ is above the diagonal $\Delta$ and the rectangle $U_{\text{left}},U_{\text{left}}^{C}]\times[V_{\text{left}},V_{\text{left}}^{C}]$ does not intersect $\Delta$. Likewise, w.h.p. the point $Z_{\text{up}}$ is above the diagonal and the rectangle $U_{\text{up}}^{C},U_{\text{up}}]\times[V_{\text{up}}^{C},V_{\text{up}}]$ does not intersect $\Delta$. Under those events, the records of $\mathcal{P}$ are all contained in
@@
This zone is shown on the left-hand side of Figure˜6. Note that the rectangles $U_{\text{left}},U_{\text{left}}^{C}]\times[V_{\text{left}},V_{\text{left}}^{C}]$ and $U_{\text{up}}^{C},U_{\text{up}}]\times[V_{\text{up}}^{C},V_{\text{up}}]$ may be flat, i.e. the identities $U_{\text{left}}=U_{\text{left}}^{C}$ and $V_{\text{up}}^{C}=V_{\text{up}}$ may occur with non-negligible probability.
Invertly, all points of $\mathcal{P}_{\Delta}$ in $0,U_{\text{left}}]^{2}\cup[V_{\text{up}},1]^{2}$, as well as the ones which are high records in $\mathcal{P}\cap C$, are high records in $\mathcal{P}$. From the previous two observations we deduce the following inequality w.h.p. as $n\to\infty$:
@@
where $S_{n}:=\left|\mathcal{P}_{\Delta}\cap[0,U_{\text{left}}]^{2}\right|+\left|\mathcal{P}_{\Delta}\cap[V_{\text{up}},1]^{2}\right|$. For any $M>0$, we may apply ˜3.2 to the rectangle $0,M/\check{n}]\times[0,1]$ to obtain $\left|\check{\mathcal{P}}\cap\left([0,M/\check{n}]\times[0,1]\right)\right|\leq 2M$ w.h.p. as $n\to\infty$. Since $\check{n}U_{\text{left}}^{C}$ is tight by ˜6.1, it holds that $U_{\text{left}}^{C}\leq M/\check{n}$ with probability tending to $1$ as $M\to\infty$, uniformly in $n$. This implies that the random variable $\left|\check{\mathcal{P}}\cap\left([U_{\text{left}},U_{\text{left}}^{C}]\times[V_{\text{left}},V_{\text{left}}^{C}]\right)\right|$ is a $\mathcal{O}_{\mathbf{P}}(1)$. With the same reasoning applied to $1-V_{\text{up}}^{C}$ and the corresponding rectangle, we deduce:
@@
Now it remains to analyze both terms in the sum. According to ˜3.3 and ˜3.2, $\check{\mathcal{P}}\cap C$ is a family of i.i.d. uniform points in $C$, with size $\check{n}/{4}+o_{\mathbf{P}}\left(\check{n}\right)$. We can thus apply Theorem˜2.9 to deduce
@@
in distribution. Let us turn to the fixed points. $\check{\mathcal{P}}$ and $\mathcal{P}_{\Delta}$ are independent, so $(U_{\text{left}},V_{\text{up}})$ and $\mathcal{P}_{\Delta}$ are as well. Therefore, conditionally given $(U_{\text{left}},V_{\text{up}})$, $S_{n}$ follows a ${\rm Binom}\left(t_{1}^{(n)},U_{\text{left}}\!+\!1\!-\!V_{\text{up}}\right)$ law. Recall from ˜6.1 that
@@
in distribution, where ${\rm Exp}\left(1\right)*{\rm Exp}\left(1\right)$ denotes the law of the sum of two independent ${\rm Exp}\left(1\right)$ random variables. If ${t_{1}^{(n)}}/{\check{n}}=\mathcal{O}(1)$ as $n\to\infty$ then $S_{n}=\mathcal{O}_{\mathbf{P}}(1)$ (use e.g. the Bienaymé–-Chebyshev inequality), whereas if ${t_{1}^{(n)}}/{\check{n}}\to\infty$ as $n\to\infty$ then
@@
in distribution. Now we can use Equations˜29, 30 and 31 to identify three regimes in the asymptotics of $\mathrm{rec}^{\uparrow}\left(\mathcal{P}\right)$.
@@
Assume $\frac{\check{n}}{n/\sqrt{\log n}}\to\alpha\in(0,\infty)$. Then $\log\check{n}\sim\log n$ as $n\to\infty$, and since $t_{1}^{(n)}=n-\check{n}$, we deduce $\frac{t_{1}^{(n)}}{\check{n}\sqrt{\log\left(\check{n}\right)}}\to 1/\alpha$. Thus we need to understand the interplay between the limit laws of Equations˜30 and 31. To do this, notice that conditionally given $\left(U_{\text{left}},V_{\text{up}}\right)$, the variables $S_{n}$ and $\check{\mathcal{P}}\cap C$ are independent. Recall that under this conditioning, $S_{n}$ follows a ${\rm Binom}\left(t_{1}^{(n)},U_{\text{left}}\!+\!1\!-\!V_{\text{up}}\right)$ law. According to ˜3.4, the conditional law of ${\rm Perm}\left(\check{\mathcal{P}}\cap C\right)$ given $\left(U_{\text{left}},V_{\text{up}}\right)=\left(U_{(1)},U_{(n)}\right)$ and $\left|\check{\mathcal{P}}\cap C\right|$ is uniform. From this property we deduce independence of the limit laws in Equations˜31 and 30, whence:
@@ 6.2 Low records
By ˜3.2, $\left|\mathcal{P}\cap B_{n}\right|$ is distributed like a sum of three ${\rm Binom}\left(\lfloor n/3\rfloor\pm 1,\left(1-\epsilon/\sqrt{n}\right)\big.\epsilon/\sqrt{n}\right)$, and likewise for $B_{n}^{*}$. The first claim readily follows from standard concentration inequalities. Then w.h.p. as $n\to\infty$, the sets $\mathcal{P}\cap B_{n}$ and $\mathcal{P}\cap B_{n}^{*}$ are non-empty and $Z_{\text{down}}^{B}$, $Z_{\text{left}}^{B^{*}}$ are well-defined. According to ˜3.3, conditionally given $\left|\mathcal{P}\cap B_{n}\right|$, the set $\mathcal{P}\cap B_{n}$ consists of i.i.d. uniform points in $B_{n}$. We can thus write, for any $s>0$ and large enough $n$, conditionally given $\left|\mathcal{P}\cap B_{n}\right|$:
@@
Define $C_{n}:=\left[0,\epsilon/\sqrt{\check{n}}\right]^{2}$. Then $\left|\check{\mathcal{P}}\cap C_{n}\right|$ is distributed like a sum of three ${\rm Binom}\left(\lfloor\check{n}/3\rfloor\pm 1,\epsilon^{2}/\check{n}\right)$ variables by ˜3.2, and $\left|\mathcal{P}_{\Delta}\cap C_{n}\right|$ follows a ${\rm Binom}\left(t_{1}^{(n)},\epsilon/\sqrt{\check{n}}\right)$ law. Therefore, there exists $\delta(\epsilon)=\underset{\epsilon\to 0}{o}\left(1\right)$, independent of $n$, such that:
@@
Define $B_{n}:=\left[0,\epsilon/\sqrt{\check{n}}\right]\times\left[\epsilon/\sqrt{\check{n}},1\right]$ and $B_{n}^{*}:=\left[\epsilon/\sqrt{\check{n}},1\right]\times\left[0,\epsilon/\sqrt{\check{n}}\right]$. According to ˜6.2 we can define w.h.p. $Z_{\text{down}}^{B}=\left(U_{\text{down}}^{B},V_{\text{down}}^{B}\right)$ as the downmost point in $\check{\mathcal{P}}\cap B_{n}$ and $Z_{\text{left}}^{B^{*}}=\left(U_{\text{left}}^{B^{*}},V_{\text{left}}^{B^{*}}\right)$ as the leftmost point in $\check{\mathcal{P}}\cap B_{n}^{*}$. See the right-hand side of Figure˜6 for a representation.
@@
$\left|\Big.\mathrm{rec}^{\downarrow}\left(\mathcal{P}\right)-\mathrm{rec}^{\downarrow}\left(\check{\mathcal{P}}\cap B_{n}\right)-\mathrm{rec}^{\downarrow}\left(\check{\mathcal{P}}\cap B_{n}^{*}\right)\right|\leq\left|\mathcal{P}\cap\left(\left[\epsilon/\sqrt{\check{n}},U_{\text{left}}^{B^{*}}\right]\times\left[\epsilon/\sqrt{\check{n}},V_{\text{down}}^{B}\right]\right)\right|.$ (33)
Using ˜6.2 and the fact that $t_{1}^{(n)}=\mathcal{O}\left(\sqrt{n}\right)$ we obtain:
@@
as $n\to\infty$, for fixed $\epsilon>0$. Thus it suffices to study the records in $B_{n}$ and $B_{n}^{*}$. We can apply ˜3.3, ˜6.2 and Theorem˜2.9 to obtain
@@
in probability, and likewise for $B_{n}^{*}$. Using Equations˜33 and 32 we deduce that for any $\eta>0$:
@@
in distribution. By ˜3.2 and since ${\check{n}}/{(t_{1}^{(n)})^{2}}\to 0$, w.h.p. it holds that $\check{\mathcal{P}}\cap[0,U_{\Delta}]^{2}=\emptyset$. Define two bands $B_{\Delta}:=[0,U_{\Delta}]\times[U_{\Delta},1]$ and $B_{\Delta}^{*}:=[U_{\Delta},1]\times[0,U_{\Delta}]$. Then under the event $\left\{\check{\mathcal{P}}\cap[0,U_{\Delta}]^{2}=\emptyset\right\}$ we may write:
@@
Using ˜3.2 and ˜3.3 conditionally given $U_{\Delta}$, $\check{\mathcal{P}}\cap B_{\Delta}$ is a family of i.i.d. uniform points in $B_{\Delta}$ whose size is distributed like a sum of three ${\rm Binom}\left(\lfloor\check{n}/3\rfloor\pm 1,U_{\Delta}(1-U_{\Delta})\right)$ variables. The same holds for $B_{\Delta}^{*}$.
  • If $n=\mathcal{O}\left(t_{1}^{(n)}\right)$ then, by Equation˜34, $\left|\mathcal{P}\cap B_{\Delta}\right|$ and $\left|\mathcal{P}\cap B_{\Delta}^{*}\right|$ are both $\mathcal{O}_{\mathbf{P}}(1)$ as $n\to\infty$. Thus by Equation˜35, $\mathrm{rec}^{\downarrow}\left(\mathcal{P}\right)$ is also a $\mathcal{O}_{\mathbf{P}}(1)$.
  • If $t_{1}^{(n)}=o(n)$ then, by Equation˜34, $\frac{t_{1}^{(n)}}{n}\left|\mathcal{P}\cap B_{\Delta}\right|\underset{n\to\infty}{\longrightarrow}{\rm Exp}\left(1\right)$ in distribution. Therefore:
@@
    in probability. In particular $\left|\mathcal{P}\cap B_{\Delta}\right|\to\infty$ in probability and we can apply Theorem˜2.9 to deduce:
@@
    in probability. The same holds for $B_{\Delta}^{*}$, and the previous two equations along with Equation˜35 yield:
@@
This concludes the proof of Theorem˜2.11. ∎
Proof of ˜2.12.
We use the same notation as in the proof of Theorem˜2.11. Recall that under $E_{n}$:
@@
Suppose that $t_{1}^{(n)}+2t_{2}^{(n)}=n$, i.e. that $\tau_{n}$ is a random involution with $t_{1}^{(n)}$ fixed points. Then the sets $\check{\mathcal{P}}\cap B_{n}$ and $\check{\mathcal{P}}\cap B_{n}^{*}$ are a.s. symmetric with respect to the diagonal. Therefore, ${\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}\right)$ and ${\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}^{*}\right)$ are a.s. inverse of each other, and $\mathrm{rec}^{\downarrow}\left({\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}\right)\right)=\mathrm{rec}^{\downarrow}\left({\rm Perm}\left(\check{\mathcal{P}}\cap B_{n}^{*}\right)\right)$ almost surely. Using Equations˜32, 36 and 37 this yields:
@@
Suppose that $t_{1}^{(n)}+3t_{3}^{(n)}=n$, i.e. that $\tau_{n}$ is a random product of $3$-cycles. We claim the following: conditionally on the event $E_{n}$ and given the family $\mathcal{P}\cap B_{n}$, the permutation ${\rm Perm}\left(\mathcal{P}\cap B_{n}^{*}\right)$ has size $\left|\mathcal{P}\cap B_{n}\right|$ and is uniformly random. Let us explain why.
@@
This holds for each triple $U_{3,k}^{1},U_{3,k}^{2},U_{3,k}^{3}$ in the construction of $\mathcal{P}$. Under the event $E_{n}$, only two possibilities remain: either $Z_{3,k}^{1},Z_{3,k}^{2},Z_{3,k}^{3}\in D_{n}$, or exactly one point is in $B_{n}$ and another one is in $B_{n}^{*}$, and they share a coordinate. The conditional law of $\mathcal{P}\cap B_{n}^{*}$ under $E_{n}$ and $\mathcal{P}\cap B_{n}$ can thus be described as follows: write $\mathcal{P}\cap B_{n}=\{(U_{i},V_{i})\}$; let $(W_{i})$ be i.i.d. ${\rm Unif}\left([\epsilon_{n},1]\right)$ variables; and set $\mathcal{P}\cap B_{n}^{*}=\{(W_{i},U_{i})\}$. One can easily see that for any distinct $u_{1},\dots,u_{r}$, if $W_{1},\dots,W_{r}$ are i.i.d. ${\rm Unif}\left([\epsilon_{n},1]\right)$ variables then ${\rm Perm}\left((W_{1},u_{1}),\dots,(W_{r},u_{r})\right)$ is uniformly random (e.g. switching $W_{i}$ with $W_{j}$ shows that this is the Haar measure on the symmetric group). This proves that conditionally on the event $E_{n}$ and given the family $\mathcal{P}\cap B_{n}$, the permutation ${\rm Perm}\left(\mathcal{P}\cap B_{n}^{*}\right)$ has size $\left|\mathcal{P}\cap B_{n}\right|$ and is uniformly random.
This claim implies that conditionally on the event $E_{n}$ and given the number $\left|\mathcal{P}\cap B_{n}\right|$, the random permutations ${\rm Perm}\left(\mathcal{P}\cap B_{n}\right)$ and ${\rm Perm}\left(\mathcal{P}\cap B_{n}^{*}\right)$ have size $\left|\mathcal{P}\cap B_{n}\right|$ and are independent uniform. Using Equations˜32, 36 and 37 we deduce: